Evaluate the limit as x turns to 3

Lim √(x+1) - 2 / x-3



1
Expert's answer
2021-09-13T00:05:08-0400

"\lim_{x\to3} \frac{\sqrt{x+1}-2}{x-3}\\
=\frac{0}{0}\\
\text{Therefore using l'hopital's rule we differentiate the numerator and denominator}\\
\lim_{x\to 3}(\frac{1}{2}(x+1)^{-\frac{1}{2}}=\frac{1}{4}"